OpenStudy (anonymous):

can someone help me solve this 3x*3x+20x=7 by factoring

OpenStudy (anonymous):

9x^2+20x-7=0

OpenStudy (anonymous):

\[\frac{ -10}{9 }\pm \frac{ \sqrt{163} }{ 9 }\]